Time Value of Money

 Let us take APR = 8 %. Now calculate the EAR for continuous compounding. 
Answer #1

Annual Percentage Rate, APR = 8% = 0.08

Formula: Effective Annual Rate, EAR = [1+(APR/m)]m - 1

where, m = number of months

∴ EAR = [1+(0.08/m)]m - 1

=> EAR = [1+(0.08/12)]12 - 1 [∵1 year = 12 months]

=> EAR = 1.083 - 1 = 0.083

Therefore, EAR = 8.3%


answered by: Zahidul Hossain